# Structured Reinforcement Learning for Combinatorial Decision-Making

### v1

@@ Discussion
This robustness comes at a computational cost. SRL requires about 30 minutes of training per environment, compared to shorter runtimes for SIL and PPO. The difference arises from CO-layer usage: PPO invokes it only twice per update, versus 20$\times$ in SIL and up to 61$\times$ in SRL. Runtime also varies with CO-layer complexity—e.g., GSPP has a simple layer, leading to similar runtimes, while the more complex layer in DVSP increases SRL’s runtime. The DAP runtime is further impacted by an expensive simulation and the use of two Q-networks. Although early stopping could mitigate this, the findings highlight a core limitation: SRL’s computational cost scales with CO-layer complexity.
Overall, the observations for our dynamic experiments align with our static experiments (Appendix G). SRL consistently matches SIL in performance and convergence, while PPO underperforms across all metrics. Runtimes follow the same pattern, increasing with CO-layer complexity.
In summary, our results yield three takeaways for real-world deployments: (i) Unstructured RL lacks the stability required for practical use. (ii) SIL is limited to settings with simple dynamics and access to expert demonstrations. (iii) Given the expected model and layer complexity in real-world settings, SRL offers a scalable and effective alternative solution approach at the price of a (reasonable) computational overhead when comparing it to SIL.
@@ 4 Conclusion
In this paper, we address combinatorial MDPs (C-MDPs), which present substantial challenges to current RL algorithms, despite being common in many industrial applications. Utilizing the framework of COaML-pipelines, we propose Structured Reinforcement Learning (SRL), a primal-dual algorithm using Fenchel-Young losses to train COaML-pipelines in an end-to-end fashion, thereby learning policies for C-MDPs using collected experience only. We compare SRL to Structured Imitation Learning (SIL) and unstructured RL in three static and three dynamic environments, representing typical industrial problem settings with combinatorial action spaces. The performance of SRL is competitive to SIL in the static environments and up to 78% better in the dynamic environments. SRL is consistently outperforms unstructured RL by up to 92%, additionally being more stable and converging quicker, at the cost of higher computational effort.
@@
We thank the BAIS research group at TUM for valuable comments and discussions. The work of Heiko Hoppe is supported by the Munich Data Science Institute with a Linde/MDSI PhD Fellowship.
Appendix A Problem setting: Combinatorial Markov Decision Processes
Consider a Markov Desicion Process with states $s\in\mathcal{S}$, actions $a\in\mathcal{A}(s)\subset\mathbb{R}^{d(s)}$ embedded in a vector space, rewards $r$, and transition probabilities $\mathbb{P}(s^{\prime},r|s,a)$ with next state $s^{\prime}$. In the case of combinatorial MDPs, $\mathcal{A}(s)$ is the set of feasible solutions of a combinatorial optimization problem. We introduce the moment polytope ${\mathcal{C}(s):=\operatorname{conv}\big{(}\mathcal{A}(s)\big{)}}$, which is useful when convexifying $\mathcal{A}(s)$ using regularization functions for end-to-end differentiation. In the literature on stochastic optimization [Bertsekas and Shreve, 1996], we often introduce a latent noise variable $\xi\in\Xi$ with probability $p(\xi\mid s,a)$, such that the transition to $(s^{\prime},r)$ given $(s,a,\xi)$ is deterministic. We distinguish exogenous noises, where the distribution of $\xi$ does not depend on $a$, and endogenous noises, where the distribution of $\xi$ depends on $a$.
$\bar{\pi}\in\operatorname*{argmax}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}\>r_{t}\right],$
$Q^{\pi}(s_{t},a_{t})=\mathbb{E}\left[r_{t}\right]+\gamma\>\mathbb{E}\left[\max_{\tilde{a}_{t+1}}Q^{\pi}(s_{t+1},\tilde{a}_{t+1})\right],$
with expectations with respect to $\mathbb{P}(s^{\prime},r\mid s,a)$.
In the context of RL, we usually perform three approximations: first, $\mathbb{P}(s^{\prime},r\mid s,a)$ is only exploited by sampling transitions from the environment. Second, we approximate $Q^{\pi}(s_{t+1},\tilde{a}_{t+1})$ using critic neural networks $\psi_{\beta}$ parametrized by weights $\beta$. Third, we approximate $\pi$ by sampling actions from an actor $a\sim\pi_{w}(\cdot\mid s)$, which is often a neural network $\varphi_{w}$ parametrized by weights $w$. In our case, the actor is a COaML-pipeline with statistical model $\varphi_{w}$ and CO-layer $f$.
Appendix B Combinatorial Optimization-augmented Machine Learning pipelines
@@
$a=f(\theta,s)\in\operatorname*{argmax}_{\tilde{a}\in\mathcal{A}(s)}:\langle\theta,\tilde{a}\rangle.$
@@
As a motivating example, consider the Warcraft Shortest Paths Problem [Vlastelica et al., 2020], illustrated in Figure 6 and detailed in Appendix F. In this setting, the state $s$ is given by a map image with dimensions $96\times 96\times 3$ pixels. The task is to find the cost-minimizing path from the top-left to the bottom-right corner of the map, which corresponds to the action $a$. The path cost is determined by the terrain the agent traverses, with terrain types encoded by specific color codes. To solve this task, the state must first be encoded into a structured representation. We employ a convolutional neural network to estimate scores for each cell in a $12\times 12$ grid. These scores $\theta$ are then used by Dijkstra’s algorithm to compute the path that minimizes the cumulative cell costs. Notably, neither the convolutional network nor Dijkstra’s algorithm alone can solve the raw WSPP map image. However, the COaML-pipeline efficiently learns to find cost-minimal paths through end-to-end training, combining the strengths of both components.
Figure 6: ML-CO-pipeline for the Warcraft Shortest Paths Problem: a convolutional neural network estimates cell scores based on image pixels. The CO-layer applies Dijkstra’s algorithm on the scores to create a path between the top left and the bottom right corner.
Appendix C Fenchel-Young losses
Given the actor policy illustrated by Figure 1, and target actions to imitate $\bar{a}$, one could consider a classic primal-primal loss to learn the actor model, such as the $\ell^{2}$ norm. The learning problem would be of the form
$\min_{w}||f\big{(}\varphi_{w}(s),s\big{)}-\bar{a}||_{2}^{2}.$ (7)
Nonetheless, several issues arise in that case. This learning problem involves the CO-layer $f$, which is piecewise constant as explained in Section 2.1. We could smooth the latter, and rely on score function estimators to optimize Problem (7) via stochastic gradient descent. Nonetheless, the resulting optimization problem is non-convex in the output of the statistical model $\theta=\varphi_{w}(s)$, and the score function estimator is known to suffer from high variance [Mohamed et al., 2020]. Those limitations make learning tough in practice, with stochastic gradient descent unable to minimize the loss properly. To mitigate these osbtacles, we consider the Fenchel-Young loss [Blondel et al., 2020]. Given a regularization function $\Omega:\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{+\infty\}$ and its Fenchel conjugate $\Omega^{*}$, the Fenchel-Young loss $\mathcal{L}_{\Omega}(\theta;\bar{a})$ generated by $\Omega$ is defined over $\operatorname{dom}(\Omega^{*})\times\operatorname{dom}(\Omega)$ as
$\mathcal{L}_{\Omega}(\theta;\bar{a}):=\Omega^{*}(\theta)+\Omega(\bar{a})-\langle\theta|\bar{a}\rangle=\sup_{a\in\operatorname{dom}(\Omega)}\big{(}\langle\theta|a\rangle-\Omega(a)\big{)}-\big{(}\langle\theta|\bar{a}\rangle-\Omega(\bar{a})\big{)}.$
For a given $\theta\in\operatorname{dom}(\Omega^{*})$, we introduce the regularized prediction as $\sup_{a\in\operatorname{dom}(\Omega)}\langle\theta|a\rangle-\Omega(a)$. The Fenchel-Young loss measures the non-optimality of $\bar{a}\in\operatorname{dom}(\Omega)$ as a solution of the regularized prediction problem. It is nonnegative and convex in $\theta$. If in addition $\Omega$ is proper, convex, and lower semi-continuous, $\mathcal{L}_{\Omega}$ reaches zero if and only if $\bar{a}$ is a solution of the regularized prediction problem. Over the last few years, Fenchel-Young losses have become the main approach for supervised training of structured policies, as they lead to a convex and tractable learning problem.
Appendix D Proofs
@@
D.1 Two regularizations: negentropy and sparse perturbation.
@@
D.2 Extension on the boundary
The expectation in the right-hand side of Equation (9) naturally extends $\nabla\Omega_{\varepsilon,\Delta}^{*}$ in ({}$\big{(}\mathbb{R}\cup\{-\infty\}\big{)}^{\mathcal{A}}$. Indeed, for any ${\gamma\in\big{(}\mathbb{R}\cup\{-\infty\}\big{)}^{\mathcal{A}}}$, and irrespective of the perturbation $Z\in\mathbb{R}^{d}$, any action $a$ satisfying $\gamma_{a}=-\infty$ will never appear in the $\operatorname*{argmax}$. This observation allows us to formally define the effective support of $\gamma$ as $\hat{\mathcal{A}}(\gamma)=\{a\in\mathcal{A}\mid\gamma_{a}>-\infty\}$, we get
@@ D.2 Extension on the boundary
where $\hat{\gamma}\in\mathbb{R}^{\hat{\mathcal{A}}(\gamma)}$ is the vector of finite components of $\gamma$, and $\nabla\Omega_{\Delta^{\hat{\mathcal{A}}(\gamma)}}^{*}(\hat{\gamma})_{a}$ is the component indexed by $a$ of the vector $\nabla\Omega_{\Delta^{\hat{\mathcal{A}}(\gamma)}}^{*}(\hat{\gamma})$ which belongs to $\Delta^{\hat{\mathcal{A}}(\gamma)}$. In the other way round, we extend $\partial\Omega_{\Delta}$ at the relative boundary of the simplex $\Delta^{\mathcal{A}}$ as follows. Let $q\in\Delta^{\mathcal{A}}$ be a sparse distribution (with some null components). In a similar way, we introduce the set $\hat{\mathcal{A}}(q)=\{a\in\mathcal{A}\mid q_{a}>0\}$, and define
$\partial\Omega_{\Delta}(q)\ni\gamma=\begin{cases}-\infty&\text{for }a\in\mathcal{A}\setminus\hat{\mathcal{A}}(q)\\ \hat{\gamma}_{a},\hat{\gamma}\in\partial\Omega_{\Delta^{\hat{\mathcal{A}}(q)}}(\hat{q})&\text{for }a\in\hat{\mathcal{A}}(q),\end{cases}$ (11)
@@ D.3 Proof of Proposition 2
where $\ln(0)$ is taken equal to $\infty$- ∞.
In step (6d), using again the extension of $\nabla\Omega_{\Delta}^{*}$ in ({}$\big{(}\mathbb{R}\cup\{-\infty\}\big{)}^{\mathcal{A}}$, Equation (8), and defining the normalization constant $\mathcal{Z}_{m,\beta}:=\sum_{a\in\mathcal{A}}\exp\big{(}\ln(k_{a})+\frac{1}{\tau}Q_{\psi_{\beta}}(a)\big{)}$,
@@ E.1.2 Hyperparameters
No. samples for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ Yes No Yes
Standard dev. $\varepsilon$ for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ Yes No Yes
E.1.3 Experiment conduction
@@
We separate all instances into a train, validation, and test dataset. We create the training dataset for SIL by applying the expert policy to the training instances and storing the solutions. To tune the hyperparameters, we use the same random model initialization for SIL, PPO, and SRL. We tune the number of episodes and the number of iterations per episode using PPO, as it is typically the most constrained in terms of iterations and requires the largest number of episodes due to its on-policy, unstructured nature. This setup favors the baselines—particularly PPO—since SRL often converges more quickly but incurs higher computational cost per episode. As a result, PPO holds a natural advantage in runtime comparisons.
@@
We then use the same number of episodes and iterations to train both PPO and SRL, and adjust the number of epochs for SIL to ensure that all methods perform approximately the same number of update steps overall. In most cases, this results in an equal number of episodes and epochs. For the DAP and GSPP, however, we reduce the number of epochs to account for the large size of the training dataset. We tune the exploration standard deviation $\sigma_{f}$, the perturbation standard deviation $\sigma_{b}$, the temperature parameter $\tau$, and the learning rate—typically shared between actor and critic, or set slightly higher for the critic—using a grid search for each algorithm. Each grid search involves between 3 and 30 training runs. All other hyperparameters do not require detailed tuning.
@@
Once the optimal hyperparameters are identified, we run each algorithm with ten randomly initialized actor (and, where applicable, critic) models. After each episode or epoch, the actor is evaluated on the training and validation datasets—or a subset thereof to improve efficiency. We save the actor model whenever it achieves the best performance observed so far. At the end of training, the best-performing actor model is restored and used for final evaluation on the training and test datasets.
E.1.4 Results metrics
@@
To compare performance, we run the best saved model of each algorithm after each run with different random model initializations on the train and the test dataset. We calculate the mean over the ten models per instance of the train and test dataset, using these mean per-instance rewards in the results boxplots. To compare convergence speed, we store the validation rewards over the course of training and calculate the mean across the ten runs per algorithm and environment per episode. In the lineplots, we display the highest mean validation reward achieved by the model so far for each training episode. We report the mean of the standard deviations across the validation and final test rewards of the ten runs in the tables. We further measure the time to run an algorithm in minutes, reporting that number in the tables as well. Finally, we calculate the overall mean reward of the final tests on the train and test dataset per algorithm and environment and display it in Appendix G.
E.2 Algorithm specification
@@
E.2.1 Structured Imitation Learning
@@
E.2.2 Proximal Policy Optimization
@@ E.2.3 Benchmark reasoning
SAC requires the actor to output both the mean and the standard deviation of a continuous action distribution in order to perform entropy regularization. This would necessitate neural networks with two outputs—$\theta$ and $\sigma_{f}$. In contrast, PPO allows us to manually set $\sigma_{f}$, avoiding the need for a separate network head or specialized actor architecture. This ensures consistency across algorithms and avoids introducing additional sources of divergence.
@@
Appendix F Environment specification
@@
F.1 Warcraft Shortest Paths Problem
@@
No. samples for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 20 20
Standard dev. $\varepsilon$ for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 0.05 0.05
@@
F.2 Single Machine Scheduling Problem
@@
Temperature param. $\tau$ (incl. schedule) 1.0
No. samples for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 20 20
Standard dev. $\varepsilon$ for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 1.0 1.0
@@
F.3 Stochastic Vehicle Scheduling Problem
@@
No. samples for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 20 20
Standard dev. $\varepsilon$ for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 1.0 1.0
@@
A feasible decision for this problem is therefore a set of disjoint $s-t$ paths such that all tasks are covered.
@@
F.4 Dynamic Vehicle Scheduling Problem
@@
No. samples for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 20 20
Standard dev. $\varepsilon$ for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 0.01 0.01
@@
F.5 Dynamic Assortment Problem
@@
No. samples for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 20 20
Standard dev. $\varepsilon$ for $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ 1.0 1.0
@@
F.6 Gridworld Shortest Paths Problem

### v2

@@ Discussion
[…]
Overall, the observations for our dynamic experiments align with our static experiments (Appendix E). SRL consistently matches SIL in performance and convergence, while PPO underperforms across all metrics. Runtimes follow the same pattern, increasing with CO-layer complexity.
In summary, our results yield three takeaways for real-world deployments: i) unstructured RL lacks the stability required for practical use. ii) SIL is limited to settings with simple dynamics and access to expert demonstrations. iii) Given the expected model and layer complexity in real-world settings, SRL offers a scalable and effective alternative solution approach at the price of a (reasonable) computational overhead when comparing it to SIL.
@@ 4 Conclusion
In this paper, we address combinatorial MDPs (C-MDPs), which present substantial challenges to current RL algorithms, despite being common in many industrial applications. Utilizing the framework of COAML-pipelines, we propose Structured Reinforcement Learning (SRL), a primal-dual algorithm using Fenchel-Young losses to train COAML-pipelines in an end-to-end fashion, thereby learning policies for C-MDPs using collected experience only. We compare SRL to Structured Imitation Learning (SIL) and unstructured RL in three static and three dynamic environments, representing typical industrial problem settings with combinatorial action spaces. The performance of SRL is competitive to SIL in the static environments and up to 78% better in the dynamic environments. SRL consistently outperforms unstructured RL by up to 92%, additionally being more stable and converging quicker, at the cost of higher computational effort.
@@
We thank the BAIS research group at TUM for valuable comments and discussions. The work of Heiko Hoppe was supported by the Munich Data Science Institute with a Linde/MDSI PhD Fellowship.
Appendix A Combinatorial Optimization-augmented Machine Learning pipelines
Figure 6: COAML-pipeline for the Warcraft Shortest Paths Problem: a convolutional neural network estimates cell scores based on image pixels. The CO-layer applies Dijkstra’s algorithm on the scores to create a path between the top left and the bottom right corner.
@@ Appendix A Combinatorial Optimization-augmented Machine Learning pipelines
$f(\theta,s):a=\operatorname*{argmax}_{\tilde{a}\in\mathcal{A}(s)}:\langle\theta,\tilde{a}\rangle$.
@@
As a motivating example, consider the Warcraft Shortest Paths Problem [Vlastelica et al., 2020], illustrated in Figure 6 and detailed in Appendix D. In this setting, the state $s$ is given by a map image with dimensions $96\times 96\times 3$ pixels. The task is to find the cost-minimizing path from the top-left to the bottom-right corner of the map, which corresponds to the action $a$. The path cost is determined by the terrain the agent traverses, with terrain types encoded by specific color codes. To solve this task, the state must first be encoded into a structured representation. We employ a convolutional neural network to estimate scores for each cell in a $12\times 12$ grid. These scores $\theta$ are then used by Dijkstra’s algorithm to compute the path that minimizes the cumulative cell costs. Notably, neither the convolutional network nor Dijkstra’s algorithm alone can solve the raw WSPP map image. However, the COAML-pipeline efficiently learns to find cost-minimal paths through end-to-end training, combining the strengths of both components.
Appendix B Proofs
@@
B.1 Two regularizations: negentropy and sparse perturbation.
@@
B.2 Extension on the boundary
The expectation in the right-hand side of Equation (10) naturally extends $\nabla\Omega_{\varepsilon,\Delta}^{*}$ in $\big(\mathbb{R}\cup\{-\infty\}\big)^{\mathcal{A}}$. Indeed, for any $\gamma\in\big(\mathbb{R}\cup\{-\infty\}\big)^{\mathcal{A}}$}, and irrespective of the perturbation $Z\in\mathbb{R}^{d}$, any action $a$ satisfying $\gamma_{a}=$-$\infty$ will never appear in the $\operatorname*{argmax}$. This observation allows us to formally define the effective support of $\gamma$ as $\hat{\mathcal{A}}(\gamma)=\{a\in\mathcal{A}\mid\gamma_{a}>-\infty\}$, we get
@@ B.2 Extension on the boundary
[…]
@@ B.3 Proof of Proposition 2
where $\ln(0)$ is taken equal to -$\infty$.
In step (8d), using again the extension of $\nabla\Omega_{\Delta}^{*}$ in $\big(\mathbb{R}\cup\{-\infty\}\big)^{\mathcal{A}}$, Equation (9), and defining the normalization constant $\mathcal{Z}_{m,\beta}:=\sum_{a\in\mathcal{A}}\exp\big(\ln(k_{a})+\frac{1}{\tau}Q_{\psi_{\beta}}(a)\big)$,
@@ C.1.2 Hyperparameters
[…]
C.1.3 Experiment conduction
@@
We separate all instances into a train, validation, and test dataset. We create the training dataset for SIL by applying the expert policy to the training instances and storing the solutions. To tune the hyperparameters, we use the same random model initialization for SIL, PPO, and SRL. We tune the number of episodes and the number of iterations per episode using PPO, as it is typically the most constrained in terms of iterations and requires the largest number of episodes due to its on-policy, unstructured nature. This setup favors the baselines – particularly PPO – since SRL often converges more quickly but incurs higher computational cost per episode. As a result, PPO holds a natural advantage in runtime comparisons.
@@
We then use the same number of episodes and iterations to train both PPO and SRL, and adjust the number of epochs for SIL to ensure that all methods perform approximately the same number of update steps overall. In most cases, this results in an equal number of episodes and epochs. For the DAP and GSPP, however, we reduce the number of epochs to account for the large size of the training dataset. We tune the exploration standard deviation $\sigma_{f}$, the perturbation standard deviation $\sigma_{b}$, the temperature parameter $\tau$, and the learning rate – typically shared between actor and critic, or set slightly higher for the critic – using a grid search for each algorithm. Each grid search involves between 3 and 30 training runs. All other hyperparameters do not require detailed tuning.
@@
Once the optimal hyperparameters are identified, we run each algorithm with ten randomly initialized actor (and, where applicable, critic) models. After each episode or epoch, the actor is evaluated on the training and validation datasets – or a subset thereof to improve efficiency. We save the actor model whenever it achieves the best performance observed so far. At the end of training, the best-performing actor model is restored and used for final evaluation on the training and test datasets.
C.1.4 Results metrics
@@
To compare performance, we run the best saved model of each algorithm after each run with different random model initializations on the train and the test dataset. We calculate the mean over the ten models per instance of the train and test dataset, using these mean per-instance rewards in the results boxplots. To compare convergence speed, we store the validation rewards over the course of training and calculate the mean across the ten runs per algorithm and environment per episode. In the lineplots, we display the highest mean validation reward achieved by the model so far for each training episode. We report the mean of the standard deviations across the validation and final test rewards of the ten runs in the tables. We further measure the time to run an algorithm in minutes, reporting that number in the tables as well. Finally, we calculate the overall mean reward of the final tests on the train and test dataset per algorithm and environment and display it in Appendix E.
C.2 Algorithm specification
@@
C.2.1 Structured Imitation Learning
@@
C.2.2 Proximal Policy Optimization
@@ C.2.3 Benchmark reasoning
[…]
@@
Appendix D Environment specification
@@
D.1 Warcraft Shortest Paths Problem
@@
[…]
@@
D.2 Single Machine Scheduling Problem
@@
Standard dev. $\sigma_{b}$ for $\widehat{a}$ (incl. schedule) 2.0
Temperature param. $\tau$ (incl. schedule) 1e3
[…]
@@
D.3 Stochastic Vehicle Scheduling Problem
@@
[…]
@@
A feasible decision for this problem is therefore a set of disjoint sts-t paths such that all tasks are covered.
@@
D.4 Dynamic Vehicle Scheduling Problem
@@
[…]
@@
D.5 Dynamic Assortment Problem
@@
[…]
@@
D.6 Gridworld Shortest Paths Problem